Circular No. 7971

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN SAGITTARIUS W. Liller, Viña del Mar, Chile, reports his discovery of a possible nova

(mag ≈ 8.5) on three Technical Pan films taken with an 80-mm f/1.4 camera lens (+ orange filter) on Sept. 15.110–15.124 UT. A CCD image by Liller taken on Sept. 16.007 yields broadband-V mag 8.13 and the following position: α = 18^{h}02^{m}22^{s}.07 ± 0^{s}.15, δ = −25^{o}20′30′′.4 ± 1′′.5 (equinox 2000.0); nothing appears at this location to mag 11.5 on two photographs taken on Sept. 11.12. Visual magnitude estimates (some forwarded by T. Kato, University of Kyoto, and E. Waagen, AAVSO): Sept. 16.082, 8.0 (B. King, Duluth, MN); 16.177, 8.1 (C. Puig, Santa Clara, CA); 16.303, 8.0 (J. Bedient, Honolulu, HI); 16.391, 7.9 (M. Linnolt, Honolulu, HI).

SUPERNOVAE 2002fc, 2002fd, 2002fe, 2002ff, 2002fg, 2002fh N. Yasuda, National Astronomical Observatory of Japan, on behalf of

the Supernova Cosmology Project (cf. IAUC 7763 and 7764, plus V. Prasad, G. Sainton, E. Smith, and A. Spadafora) and the Subaru high-redshift supernova search group (M. Doi, K. Nariai, M. Watanabe, Y. Urata, Y. Komiyama, N. Kashikawa, T. Yamada, M. Ouchi, T. Murayama, K. Shimasaku, H. Furusawa), reports the discovery of six supernovae found with the Subaru telescope (+ prime-focus Suprime-Cam). Nothing was visible at the respective positions to limiting magnitude i′ ∼ 26.0 on images taken on Apr. 9 (SN 2002fh), 12 (SN 2002fg), and Mar. 19 (the rest).

SN 2002 UT α_{2000} δ_{2000} i′ z Type
2002fc Apr. 9.4 14^{h}00^{m}23.^{s}29 + 5^{o}45′41′′.9 24.2 0.88 Ia
2002fd  9.5 14 03 54.08 + 4 59 49.0 23.4 0.278 Ia
2002fe  9.5 14 04 18.16 + 5 19 25.6 24.7 1.086 Ia
2002ff 9.6 14 03 39.82 + 5 46 50.6 24.0 1.1 Ia?
2002fg May 7.5 13 24 25.92 +27 44 33.9 23.7 0.79 Ia?
2002fh May 9.3 13 25 46.62 +27 34 21.1 22.4 0.33 Ic?

The supernova spectra were obtained at the Very Large Telescope on Apr. 11 and June 5; at the Gemini telescope on Apr. 16, May 11, and June 5; and at the Subaru telescope on Apr. 16. The redshifts were determined from host-galaxy emission and absorption lines, and all the supernovae were found to be within 2′′ of their host-galaxies’ centers. Photometry was also obtained with the Hubble Space Telescope during Apr. 23–July 3 for SNe 2002fc, 2002fe, 2002ff, and 2002fg.

2002 September 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green